A body on the circular orbit makes an angular displacement given by(t)= 2t²+5t+5r, if t is in seconds calculate the angular velocity at t= 2s
1
Expert's answer
2020-07-22T10:33:34-0400

By definition, the angular velocity is:


"\omega(t) = \dfrac{d\phi(t)}{dt}"

where "\phi(t) = 2t^2 + 5t + 5" is the angular displacement. Thus, have:


"\omega(t) = \dfrac{d\phi(t)}{dt} = 4t+5 \space (rad/s)"

At t= 2s:


"\omega(2s) = 4\cdot 2s + 5 = 8+5=13\space rad/s"

Answer. 13 rad/s.